# iHERO: Interactive Human-oriented Exploration and Supervision Under Scarce Communication

### v1

@@ Abstract
Exploration of unknown scenes before human entry is essential for safety and efficiency in numerous scenarios, e.g., subterranean exploration, reconnaissance, search and rescue missions. Fleets of autonomous robots are particularly suitable for this task, via concurrent exploration, multi-sensory perception and autonomous navigation. Communication however among the robots can be severely restricted to only close-range exchange via ad-hoc networks. Although some recent works have addressed the problem of collaborative exploration under restricted communication, the crucial role of the human operator has been mostly neglected. Indeed, the operator may: (i) require timely update regarding the exploration progress and fleet status; (ii) prioritize certain regions; and (iii) dynamically move within the explored area; To facilitate these requests, this work proposes an interactive human-oriented online coordination framework for collaborative exploration and supervision under scarce communication (iHERO). The robots switch smoothly and optimally among fast exploration, intermittent exchange of map and sensory data, and return to the operator for status update. It is ensured that these requests are fulfilled online interactively with a pre-specified latency. Extensive large-scale human-in-the-loop simulations and hardware experiments are performed over numerous challenging scenes, which signify its performance such as explored area and efficiency, and validate its potential applicability to real-world scenarios.
@@ I Introduction
Exploration of unknown and hazardous scenes before allowing humans inside is crucial for their safety, which can also improve the performance of subsequent tasks with an accurate model. Fleets of UAVs and UGVs have been deployed as extended eyes and arms, e.g., to explore planetary caves in Petráček et al. [17], Klaesson et al. [11]; search and rescue after earthquakes in Couceiro [5]. Many collaborative exploration strategies have been proposed along with the advances in autonomous navigation and perception, e.g., Yamauchi [23], Hussein et al. [10], Colares and Chaimowicz [4], Zhou et al. [25], Patil et al. [15]. However, they often assume an all-to-all communication among the robots, where any map observed locally by one robot is immediately available to other robots. This could be impractical in many aforementioned scenes where the communication facilities are unavailable or severely degraded. In such cases, the robots can only exchange information via ad-hoc networks in close proximity. This imposes great challenges on the fleet coordination as communication events and exploration tasks are now closely dependent, thus should be planned simultaneously.
@@ I-B Our Method
Main contribution of this work is threefold: (i) the novel problem formulation of human-oriented collaborative exploration and supervision under scarce communication, supporting three types of requests. To the best of our knowledge, this problem has not been addressed in related work; (ii) the distributed coordination strategy for both fast exploration and inter-robot or robot-operator communication, while ensuring a bounded information latency and a timely response to online requests; (iii) the numerical results from extensive simulation and hardware experiments validate its potential applicability to real-world scenarios.
@@ III-A Intermittent Communication Protocol
To ensure that the information can be propagated among the team and to the operator, the robots are required to meet and communicate via the protocol of intermittent communication during exploration. To begin with, a communication topology is defined as an undirected graph $\mathcal{G}\triangleq(\mathcal{N},\mathcal{E})$, where $\mathcal{E}\subset\mathcal{N}\times\mathcal{N}$. It satisfies two conditions: (i) $\mathcal{G}$ is connected; (ii) two robots $i,j$ can only communicate at time $t$ if $(i,j)\in\mathcal{E}$ and $\texttt{Qual}(\cdot)$ in (2) holds. Given a topology $\mathcal{G}$, the protocol is as follows: initially, all robots are in close proximity but each robot only communicates with its neighbors in $\mathcal{G}$; during each communication, two robots determine locally the “time and place” for their next communication event, and whether one robot needs to return to the operator before they meet (called a return event); then they depart and do not communicate until the next communication event. This procedure repeats indefinitely until termination, as shown in Fig. 4.
@@ 20 $\widehat{\Gamma}^{+}_{j}\leftarrow\widehat{\mathbf{p}}^{K_{j}+1}_{j}+\widehat{c}_{ij}$;
as the estimated time for robots $i$ or $j$ to reach $f$ and then return to the operator; Second, if there is a prioritized region by $Q_{1}$ request, the cost is given by $cost(f)\triangleq T_{i}^{\texttt{nav}}(f,\,p_{c})$, where $p_{c}$ is the center of the prioritized region. As proven in Lemma 2, the above procedure terminates in finite steps.
@@ V-C Proof of Theorem 1
Lemma 2 has ensured that the coordination during each pairwise communication is guaranteed to return a feasible solution. Thus, it remains to show that within the spread mode, the two conditions of Proposition 1 are satisfied. Regarding the first condition, Alg. 1 ensures that all communication events satisfy the constraint (11), meaning that the time of the next return event is bounded by $t_{r+1}\leq\mathbf{min}_{n}\{\Omega^{\texttt{h}}_{ij}[n]\}+T_{\texttt{h}}\leq\chi^{\texttt{h}}_{r}+T_{\texttt{h}}$. The second condition is ensured by the cyclic communication topology and the pairwise coordination process described above. More specifically, assume that robot $i$ returns to the operator before communicating with its succeeding robot $j$ in $\mathcal{G}_{\texttt{S}}$ during round $r$. Due to the cyclic communication topology, $\mathbf{argmin}_{n}\{T^{\texttt{h}}_{r}[n]\}=j$. Considering the strategy of choosing the robot to return as described before, the robot that returns to base in round $r+1$ is either $j$ or any other succeeding robots of $j$, which means that $T^{\texttt{h}}_{r+1}$ should contain the information of robot $i$ and $j$’s meeting event. Without loss of generality, assume that robot $j$’s succeeding robot is $k$, then it holds that $\mathbf{min}_{n}\{T^{\texttt{h}}_{r+1}[n]\}\geq t_{ij}>t^{-}_{jk}=\mathbf{min}_{n}\{T^{\texttt{h}}_{r}[n]\}$, where $t^{-}_{jk}$ is the meeting time of communication event between robot $j$ and $k$ that happens before robot $i$ returns, and $t_{ij}$ is the meeting time of $i$ and $j$ after $i$ returns, implying that the second condition is satisfied. And if the adopted topology is not a ring topology, the choice of return robot still ensures that the largest latency will be reduced via the return events before it exceeds the bound. This completes the proof. ∎

### v2

@@ Abstract
Exploration of unknown scenes before human entry is essential for safety and efficiency in numerous scenarios, e.g., subterranean exploration, reconnaissance, search and rescue missions. Fleets of autonomous robots are particularly suitable for this task, via concurrent exploration, multi-sensory perception and autonomous navigation. Communication however among the robots can be severely restricted to only close-range exchange via ad-hoc networks. Although some recent works have addressed the problem of collaborative exploration under restricted communication, the crucial role of the human operator has been mostly neglected. Indeed, the operator may: (i) require timely update regarding the exploration progress and fleet status; (ii) prioritize certain regions; and (iii) dynamically move within the explored area; To facilitate these requests, this work proposes an interactive human-oriented online coordination framework for collaborative exploration and supervision under scarce communication (iHERO). The robots switch smoothly and optimally among fast exploration, intermittent exchange of map and sensory data, and return to the operator for status update. It is ensured that these requests are fulfilled online interactively with a pre-specified latency. Extensive large-scale human-in-the-loop simulations and hardware experiments are performed over numerous challenging scenes, which signify its performance such as explored area and efficiency, and validate its potential applicability to real-world scenarios. The videos are available on https://zl-tian.github.io/iHERO/.
@@ I Introduction
Exploration of unknown and hazardous scenes before allowing humans inside is crucial for their safety, which can also improve the performance of subsequent tasks with an accurate model. Fleets of UAVs and UGVs have been deployed as extended eyes and arms, e.g., to explore planetary caves in Petráček et al. [20], Klaesson et al. [14]; search and rescue after earthquakes in Couceiro [5]. Many collaborative exploration strategies have been proposed along with the advances in autonomous navigation and perception, e.g., Yamauchi [26], Hussein et al. [12], Colares and Chaimowicz [4], Zhou et al. [28], Patil et al. [18], Guo and Dimarogonas [8]. However, they often assume an all-to-all communication among the robots, where any map observed locally by one robot is immediately available to other robots. This could be impractical in many aforementioned scenes where the communication facilities are unavailable or severely degraded. In such cases, the robots can only exchange information via ad-hoc networks in close proximity. This imposes great challenges on the fleet coordination as communication events and exploration tasks are now closely dependent, thus should be planned simultaneously.
@@ I-B Our Method
Main contribution of this work is threefold: (i) the novel problem formulation of human-oriented collaborative exploration and supervision under scarce communication, supporting three types of requests. To the best of our knowledge, this problem has not been addressed in related work; (ii) the distributed coordination strategy for both fast exploration and inter-robot-operator communication, while ensuring a timely response to online requests; (iii) the numerical validation from extensive simulation and hardware experiments for its applicability to real-world scenarios.
@@ III-A Intermittent Communication Protocol
To ensure that the information can be propagated among the team and to the operator, the robots are required to meet and communicate via the protocol of intermittent communication during exploration Guo and Zavlanos [9], Kantaros et al. [13]. To begin with, a communication topology is defined as an undirected graph $\mathcal{G}\triangleq(\mathcal{N},\mathcal{E})$, where $\mathcal{E}\subset\mathcal{N}\times\mathcal{N}$. It satisfies two conditions: (i) $\mathcal{G}$ is connected; (ii) two robots $i,j$ can only communicate at time $t$ if $(i,j)\in\mathcal{E}$ and $\texttt{Qual}(\cdot)$ in (2) holds. Given a topology $\mathcal{G}$, the protocol is as follows: initially, all robots are in close proximity but each robot only communicates with its neighbors in $\mathcal{G}$; during each communication, two robots determine locally the “time and place” for their next communication event, and whether one robot needs to return to the operator before they meet (called a return event); then they depart and do not communicate until the next communication event. This procedure repeats indefinitely until termination, as shown in Fig. 4.
@@ 20 $\widehat{\Gamma}^{+}_{j}\leftarrow\widehat{\mathbf{p}}^{K_{j}+1}_{j}+\widehat{c}_{ij}$;
as the estimated time for robots $i$ or $j$ to reach $f$ and then return to the operator; Second, if there is a prioritized region by $Q_{1}$ request, the cost is given by $cost(f)\triangleq T_{i}^{\texttt{nav}}(f,\,p_{c})$, where $p_{c}$ is the center of the prioritized region, as shown in Fig. 6. As proven in Lemma 2, the above procedure terminates in finite steps.
@@
Figure 6: Illustration of the revised plans with (right) and without (left) the prioritized region.
@@ V-C Proof of Theorem 1
Lemma 2 has ensured that the coordination during each pairwise communication is guaranteed to return a feasible solution. Thus, it remains to show that the two conditions of Proposition 1 are satisfied. Regarding the first condition, Alg. 1 ensures that all communication events satisfy the constraint (11), meaning that the time of the next return event is bounded by $t_{r+1}\leq\mathbf{min}_{n}\{\Omega^{\texttt{h}}_{ij}[n]\}+T_{\texttt{h}}\leq\chi^{\texttt{h}}_{r}+T_{\texttt{h}}$. The second condition is ensured by the cyclic communication topology and the pairwise coordination process described above. More specifically, assume that robot $i$ returns to the operator before communicating with its succeeding robot $j$ in $\mathcal{G}_{\texttt{S}}$ during round $r$. Due to the cyclic communication topology, $\mathbf{argmin}_{n}\{T^{\texttt{h}}_{r}[n]\}=j$. Considering the strategy of choosing the robot to return as described before, the robot that returns to base in round $r+1$ is either $j$ or any other succeeding robots of $j$, which means that $T^{\texttt{h}}_{r+1}$ should contain the information of robot $i$ and $j$’s meeting event. Without loss of generality, assume that robot $j$’s succeeding robot is $k$, then it holds that $\mathbf{min}_{n}\{T^{\texttt{h}}_{r+1}[n]\}\geq t_{ij}>t^{-}_{jk}=\mathbf{min}_{n}\{T^{\texttt{h}}_{r}[n]\}$, where $t^{-}_{jk}$ is the meeting time of communication event between robot $j$ and $k$ that happens before robot $i$ returns, and $t_{ij}$ is the meeting time of $i$ and $j$ after $i$ returns, implying that the second condition is satisfied. And if the adopted topology is not a ring topology, consider a robot $i$ communicating with robot $j$ in $c_{ij}$, to decide their next event $c^{\prime}_{ij}$. If $i$ returns to the operator at $t_{i}$, then equation (11) implies that $t_{i}\leq t^{\texttt{h}}_{i}(t_{i})+T_{\texttt{h}}$; while if $i$ doesn’t return, the strategy of determining return events ensures that $t-t^{\texttt{h}}_{i}(t)<T_{\texttt{h}},\forall t\in[t_{ij},t^{\prime}_{ij}]$, where $t_{ij}$ is the meeting time of $i$ and $j$ in $c_{ij}$, and $t^{\prime}_{ij}$ is the meeting time of $i$ and $j$ in $c^{\prime}_{ij}$. Thus, time constraint is satisfied locally for each robot $i$, and hence the global constraint is satisfied for the whole fleet. This completes the proof. ∎